# Explainable Hyperdimensional Computing for Balancing Privacy and Transparency in Additive Manufacturing Monitoring

### v1

@@ 3 Research Methodology
Figure 1: Hyperdimensional computing framework, including encoding, training, inference, and retraining.
@@ 3.1 Hyperdimensional Computing
Hyperdimensional computing draws significant inspiration from the architecture of the human brain, utilizing high-dimensional vectors to simulate cognitive functions. This advanced approach allows HD to emulate brain-like methodologies for processing, analyzing, and storing information across diverse cognitive tasks. The HD framework, depicted in Figure 1, is structured into four key phases: encoding, training, inference, and retraining. Each phase is meticulously designed to model different aspects of cognitive processing.
@@ 3.1.1 Encoding
The encoding phase focuses on converting feature vectors into hypervectors, establishing a high-dimensional framework. This transformation ensures that each element within a hypervector uniformly contributes to the information encoding, maintaining a balanced impact on the overall data representation. Feature vectors, denoted as $\vec{F}=\{f_{j}\}_{j=1}^{J}$, are mapped into hypervectors $\vec{H}$. This mapping involves a dot product between $\vec{F}$ and a uniquely generated vector for each dimension. Enhanced by adding a random sample from a uniformly distributed sequence $\vec{U}=\{u_{i}\}_{i=1}^{D}$, adhering to the distribution $\mathit{U}(0,2\pi)$, the operation leads to the equation $h_{i}=\cos(\vec{F}\cdot\vec{B}_{i}+u_{i})$. Here, each $\vec{B}_{i}$ represents a random vector drawn from a Gaussian distribution with a mean of zero and a standard deviation (SD) of ${\sigma}_{b}$. The vectors $\{B_{i}\}_{i=1}^{D}$, serving as the random basis, are predetermined and remain constant throughout the process. This sophisticated methodology effectively converts continuous feature vectors into continuous-valued hypervectors, making them suitable for various classification tasks, including anomaly detection [63].
@@ 3.2 Differential Privacy
Table 1: Notation and description of variables used in the DP-HD framework.
Notation Description
$\vec{F}=\{f_{j}\}_{j=1}^{J}$ Feature vector
$\vec{H}$ Hypervector
$\vec{C}_{s}$ Class hypervector assigned to class $s$
$\tilde{\vec{C}}_{s}$ Noisy class hypervector
$\mathcal{N}(0,\Delta g^{2}\sigma_{dp}^{2})$ Differential privacy noise
$\Delta g$ Sensitivity of the model
$\sigma_{dp}$ Noise level
$\epsilon$ Privacy budget
$\delta$ Privacy loss threshold
@@ 3.3 Differential Privacy-Hyperdimensional Computing
Another crucial parameter that dictates the privacy level and sets the noise level is the privacy budget. The privacy budget can be selected to predict the impact of noise on the model’s performance. In Section 3.3.1, we delve into the explainability of the DP-HD mechanism. We begin by assessing the effect of noise and training samples on the model’s performance through an examination of its memorization capability. We then define the signal-to-noise ratio (SNR), which is derived from the ratio of the signal (training samples) contribution to the model’s performance against the noise contribution. This metric enables us to forecast how noise levels influence accuracy, allowing us to set specific privacy and performance levels concurrently. Following this, we explore the decodability of DP-HD. We analyze an analytical method for extracting information from a hypervector by subtracting two secure class hypervectors generated from datasets differing by only one sample. Moreover, in Section 3.3.2, we outline an algorithm to find the optimal encoding strategy to balance privacy and performance. The encoding phase is pivotal to the model’s success, as it determines how data is mapped from feature space to hyperspace.
@@ 4.1 LPBF AM Machine Instrumentation and Setup
This section outlines the instrumentation of a commercial Laser Powder Bed Fusion (LPBF) machine, specifically the EOS M270 model situated at the National Institute of Standards and Technology (NIST). Integration involved outfitting the machine with a high-speed visible camera. The camera boasts a frame rate of 400 frames per second and a spectral response spanning 300-950 nm, facilitating real-time monitoring of the LPBF process. This monitoring entails capturing disparities in the thermal signature of the part as it undergoes layer-wise construction, enhancing quality assurance. Positioned within the build chamber’s upper right corner, the high-speed visible camera is connected via data cables routed through a custom sealed through-port. Notably, precise measurements regarding the camera’s orientation and distance were not conducted. The camera utilized features a Silicon-based array and boasts a resolution of 1.2 megapixels. Its placement inside the build chamber enables comprehensive imaging of the part’s shape and surrounding spatter pattern. Images captured by the high-speed visible camera are windowed to dimensions of 256 pixels × 256 pixels and acquired at a rate of 1000 frames per second. These images serve as representative snapshots, depicting both overhang and bulk build features. The experimental focus lies on a straightforward unsupported overhang geometry, selected by NIST researchers for its suitability in investigating meltpool behavior during overhang construction, thereby facilitating subsequent physical modeling efforts [64].
@@ 5 Experimental Results
5.1 Training Privacy
Identifying the optimal SD, denoted as ${\sigma}^{*}_{\text{b}}$, is essential for efficient encoding processes. Figure 4 reveals how SD influences the quality of encoding for optical images, illustrating the balance required between exclusive and inclusive encoding approaches. A smaller ${\sigma}_{\text{b}}$, for example, less than 0.1, leads to a high degree of similarity among hypervectors, even when the images are widely separated. This similarity often surpasses 0.5 under these conditions. On the contrary, as ${\sigma}_{\text{b}}$ increases, the similarity of hypervectors starts to correspond more closely with the actual physical distances between the images. At a ${\sigma}_{\text{b}}$ of 1, there is virtually no similarity in hypervectors for images that are significantly distanced. Such findings highlight the critical importance of selecting a ${\sigma}_{\text{b}}$ that optimally balances the feature space compression and dilution, ensuring that the hypervectors are true representations of the original feature vectors.
In DP-HD, the optimal value, ${\sigma}^{*}_{\text{b}}$, achieves both higher accuracy and a more efficient privacy budget, implying increased noise and improved privacy. For a generated dataset for a classification task, with each class’s samples distributed normally around class centers, characterized by an SD of ${\sigma}_{\text{sample}}$. These centers maintain a uniform separation distance of $\Delta C$. Figure 5 depicts how varying $\Delta C$ influences ${\sigma}^{*}_{\text{b}}$. A $\Delta C$ of 5, indicating proximity between classes, results in a ${\sigma}^{*}_{\text{b}}$ of 0.23. Increasing the separation to 10 adjusts ${\sigma}^{*}_{\text{b}}$ to 0.11, illustrating the need for a more inclusive encoding strategy to generalize across similar features, as elaborated in Section 3.1.1. Additionally, the distribution spread of each class, ${\sigma}_{\text{sample}}$, also affects ${\sigma}^{*}_{\text{b}}$. As shown in Figure 6, reducing ${\sigma}_{\text{sample}}$ leads to an increase in ${\sigma}^{*}_{\text{b}}$, necessitating precise differentiation between encoded vectors through an exclusive encoding strategy discussed in Section 3.1.1. ${\sigma}^{*}_{\text{b}}$ is 0.12, 0.23, 0.33, and 0.57 when ${\sigma}_{\text{sample}}$ is 3, 1.5, 1, and 0.5, respectively.
@@ 5.1 Training Privacy
Subsequently, we adjusted the size of the hypervector and the number of training samples, key parameters in the DP-HD framework. Figure 7 illustrates the impact on accuracy across various hypervector sizes. It becomes evident that ${\sigma}^{*}_{\text{b}}$ remains unaffected by changes in hypervector size. Similarly, as depicted in Table 2, the number of training samples also does not influence ${\sigma}^{*}_{\text{b}}$. Thus, we deduce that ${\sigma}^{*}_{\text{b}}$ is independent of these DP-HD parameters and the scale of training data. It is solely contingent upon the distribution characteristics of the feature vectors, such as the proximity between class centers and the dispersion of samples around these centers.
@@
Subsequently, we evaluated our dataset to determine the optimal ${\sigma}^{*}_{\text{b}}$ for our application. The core objective of our research was to find the best compromise between robust privacy protection and high operational efficiency. Figure 8 illustrates how the DP-HD model performs under different $\epsilon$ levels with both exclusive and inclusive encoding strategies. Our efforts focused on finding a ${\sigma}^{*}_{\text{b}}$ that achieves high accuracy, indicating superior performance, while maintaining a low $\epsilon$, which reflects stronger privacy. In previous studies of HD, ${\sigma}_{\text{b}}$ is fixed at one, which we call it BaselineHD, the accuracy remains below 15% across all $\epsilon$ values, underscoring the significance of adapting ${\sigma}_{\text{b}}$ to balance privacy and performance effectively. In DP-HD, we selected ${\sigma}^{*}_{\text{b}}$, optimizing both accuracy and privacy. At this setting, the model achieves the highest accuracy of 97.6% with the least amount of noise, corresponding to the highest $\epsilon$ of 10. As $\epsilon$ decreases, privacy increases, but at the expense of accuracy. For instance, when $\epsilon$ falls to 0.2, accuracy drops below 70%. Conversely, when $\epsilon$ exceeds 0.5, accuracy remains above 90%. It is clear that for any level of privacy, as set by choosing a specific value of $\epsilon$, the ${\sigma}^{*}_{\text{b}}$ will be changed. This happen since there is a randomness in generating the random basis vectors that effect the model’s performance. However, it is clear for all privacy $\epsilon$ values, the ${\sigma}^{*}_{\text{b}}$ is close to 0.2. Therefore we chose 0.2 as ${\sigma}^{*}_{\text{b}}$ to assign our dataset.
@@
Figure 11: Decoding optical images of LPBF process to capture the confidential information.
For the concluding segment of this section, we benchmarked our newly developed DP-HD model against traditional heavyweights in the image classification arena. This comparison includes an evaluation alongside models such as ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2. To implement DP, we utilized Opacus, a freely available open-source library designed for PyTorch that facilitates training deep learning models with DP mechanisms. Table 3 details the accuracies post-DP application, revealing that DP-HD sustains a minor accuracy reduction compared to significant drops in other models. With a stringent privacy level, represented by an $\epsilon$ value of 0.6, DP-HD maintains an accuracy of 91.04%, whereas other models plummet below 43%, showcasing DP-HD’s robustness against high noise levels as it distributes memory evenly across all dimensions of the hyperspace. As $\epsilon$ increases to 0.8, reducing the noise, DP-HD’s accuracy climbs to 93.34%, significantly surpassing other models which remain under 54%. This performance gap widens further when $\epsilon$ is adjusted to 1, with DP-HD reaching an accuracy of 94.43%, and the closest competitor, DenseNet201, achieves only 69.13%, still 25.3% lower than DP-HD. As $\epsilon$ increases further to 2, the accuracy of DP-HD is 96.30%, while the other models’ accuracies—75.22% for ResNet50, 82.48% for GoogLeNet, 58.20% for AlexNet, 77.70% for DenseNet201, and 57.20% for EfficientNet B2—illustrate a substantial disparity, with GoogLeNet being the closest yet still 13.82% less accurate than DP-HD. Even at an $\epsilon$ of 3, where privacy levels are lower and noise reduction is significant, DP-HD’s accuracy remains high at 96.66%, outperforming ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2, which exhibit accuracies of 81.18%, 83.73%, 56.27%, 86.89%, and 56.46% respectively. This demonstrates DP-HD’s superior capability to balance privacy and performance compared to other prevalent ML models in the field.

### v3

@@ 3.1 Hyperdimensional Computing
Hyperdimensional Computing (HD) draws significant inspiration from the architecture of the human brain, utilizing high-dimensional vectors to simulate cognitive functions [76]. This advanced approach allows HD to emulate brain-like methodologies for processing, analyzing, and storing information across diverse cognitive tasks. The HD framework, depicted in Figure 1, is structured into four key phases: encoding, training, inference, and retraining. Each phase is meticulously designed to model different aspects of cognitive processing.
Figure 1: Hyperdimensional computing framework, including encoding, training, inference, and retraining. In encoding, feature vectors are mapped to hypervectors, where each element, $h_{i}$, is generated by projecting a feature vector onto a random basis vector $\vec{B}_{i}$, sampled from a normal distribution with mean zero and variance $\sigma_{b}^{2}$. A uniform distribution, $\vec{U}$, is added to enhance the projection. Training aggregates hypervectors from the same class to form class hypervectors, inference compares query hypervectors for similarity, and retraining adjusts hypervectors based on misclassifications.
@@ 3.1.1 Encoding
The encoding phase transforms feature vectors into high-dimensional hypervectors, ensuring a balanced data representation by distributing information uniformly across each element. As depicted in Figure 1, a feature vector $\vec{F}=\{f_{j}\}_{j=1}^{J}$ is mapped into a hypervector $\vec{H}$ by applying a dot product with a unique basis vector for each dimension. Enhanced with a uniformly distributed random sequence $\vec{U}=\{u_{i}\}_{i=1}^{D}$ sampled from $\mathit{U}(0,2\pi)$, the result is given by $h_{i}=\cos(\vec{F}\cdot\vec{B}_{i}+u_{i})$, where $\vec{B}_{i}$ represents Gaussian random vectors with mean zero and SD ${\sigma}_{b}$. These basis vectors remain consistent throughout, enabling continuous feature-to-hypervector conversion for tasks like anomaly detection [77].
@@ 3.1.4 Retraining
Table 2: Notation and description of variables used in the differential privacy-hyperdimensional computing framework.
Notation Description
$\vec{F}=\{f_{j}\}_{j=1}^{J}$ Feature vector
$J$ Feature dimension
$\vec{H}$ Hypervector
$D$ Hypervector size
$\vec{C}_{s}$ Class hypervector assigned to class $s$
$\tilde{\vec{C}}_{s}$ Noisy class hypervector
$\vec{H}_{q}$ Query hypervector
$\sigma^{*}_{b}$ Optimal standard deviation of the random basis vectors
$\mathcal{N}(0,\Delta g^{2}\sigma_{dp}^{2})$ Differential privacy noise
$\Delta g$ Sensitivity of the model
$g$ AI model
$\sigma_{dp}$ Noise level
$\epsilon$ Privacy budget
$\delta$ Privacy loss threshold
$M$ Secure AI model
$I_{1}andI_{2}$ Two neighboring datasets differ in exactly one sample
$N$ The number of training samples
$T$ The number of training epochs
$\mu_{c}$ Average similarity of training hypervectors
$\sigma_{c}$ Standard deviation of the similarity among training hypervectors
$\hat{\vec{F}}$ Reconstructed feature vector
$\vec{\xi}$ Encoder noise
$\Sigma$ Variance of the decoder noise
@@ 3.3 Differential Privacy-Hyperdimensional Computing
Another crucial parameter that dictates the privacy level and sets the noise level is the privacy budget. The privacy budget can be selected to predict the impact of noise on the model’s performance. In Section 3.3.1, we delve into the explainability of the DP-HD mechanism. We begin by assessing the effect of noise and training samples on the model’s performance through an examination of its memorization capability. We then define the SNR, which is derived from the ratio of the signal (training samples) contribution to the model’s performance against the noise contribution. This metric enables us to forecast how noise levels influence accuracy, allowing us to set specific privacy and performance levels concurrently. Following this, we explore the decodability of DP-HD. We analyze an analytical method for extracting information from a hypervector by subtracting two secure class hypervectors generated from datasets differing by only one sample. Moreover, in Section 3.3.2, we outline an algorithm to find the optimal encoding strategy to balance privacy and performance. The encoding phase is pivotal to the model’s success, as it determines how data is mapped from feature space to hyperspace.
@@ 4.1 LPBF AM Machine Instrumentation and Setup
This section outlines the instrumentation of a commercial LPBF machine, specifically the EOS M270 model situated at the National Institute of Standards and Technology (NIST). Integration involved outfitting the machine with a high-speed visible camera. The camera boasts a frame rate of 400 frames per second and a spectral response spanning 300-950 nm, facilitating real-time monitoring of the LPBF process. This monitoring entails capturing disparities in the thermal signature of the part as it undergoes layer-wise construction, enhancing quality assurance. Positioned within the build chamber’s upper right corner, the high-speed visible camera is connected via data cables routed through a custom sealed through-port. Notably, precise measurements regarding the camera’s orientation and distance were not conducted. The camera utilized features a Silicon-based array and boasts a resolution of 1.2 megapixels. Its placement inside the build chamber enables comprehensive imaging of the part’s shape and surrounding spatter pattern. Images captured by the high-speed visible camera are windowed to dimensions of 256 pixels × 256 pixels and acquired at a rate of 1000 frames per second. These images serve as representative snapshots, depicting both overhang and bulk build features. The experimental focus lies on a straightforward unsupported overhang geometry, selected by NIST researchers for its suitability in investigating meltpool behavior during overhang construction, thereby facilitating subsequent physical modeling efforts [78].
@@ 5 Experimental Results
5.1 Training Privacy
5.1.1 Sensitivity and Uncertainty Analysis
Identifying the optimal SD, denoted as ${\sigma}^{*}_{\text{b}}$, is essential for effective encoding processes. Figure 4 shows how SD influences the effectiveness of encoding for optical images, illustrating the trade-off required between exclusive and inclusive encoding approaches. A smaller ${\sigma}_{\text{b}}$, for example, less than 0.1, leads to a significant level of similarity among hypervectors, even when the images are widely separated. This similarity often surpasses 0.5 under these conditions. On the other hand, as ${\sigma}_{\text{b}}$ increases, the similarity of hypervectors starts to correspond more closely with the actual physical distances between the images. At a ${\sigma}_{\text{b}}$ of 1, there is almost zero similarity in hypervectors for images that are significantly distanced. Such findings highlight the vital role of selecting a ${\sigma}_{\text{b}}$ that optimally balances the feature space compression and dilution, ensuring that the hypervectors are true representations of the original feature vectors.
In DP-HD, the optimal parameter, ${\sigma}^{*}_{\text{b}}$, achieves both higher accuracy and a more effective privacy budget, implying enhanced noise and improved privacy. For a synthetic dataset for a classification task, with each class’s samples distributed in a Gaussian manner around class centers, characterized by an SD of ${\sigma}_{\text{sample}}$. These centroids maintain a uniform separation distance of $\Delta C$. Figure 5 illustrates how varying $\Delta C$ affects ${\sigma}^{*}_{\text{b}}$. A $\Delta C$ of 5, indicating proximity between classes, results in a ${\sigma}^{*}_{\text{b}}$ of 0.23. Increasing the separation to 10 changes ${\sigma}^{*}_{\text{b}}$ to 0.11, demonstrating the need for a more inclusive encoding strategy to generalize across similar features, as elaborated in Section 3.1.1. Additionally, the distribution spread of each class, ${\sigma}_{\text{sample}}$, also impacts ${\sigma}^{*}_{\text{b}}$. As shown in Figure 6, reducing ${\sigma}_{\text{sample}}$ results in an increase in ${\sigma}^{*}_{\text{b}}$, necessitating precise differentiation between encoded vectors through an exclusive encoding strategy explained in Section 3.1.1. ${\sigma}^{*}_{\text{b}}$ is 0.12, 0.23, 0.33, and 0.57 when ${\sigma}_{\text{sample}}$ is 3, 1.5, 1, and 0.5, respectively.
@@ 5.1.1 Sensitivity and Uncertainty Analysis
Subsequently, we modified the size of the hypervector and the number of training samples, key factors in the DP-HD framework. Figure 7 demonstrates the impact on accuracy across various hypervector sizes. It becomes evident that ${\sigma}^{*}_{\text{b}}$ remains unaffected by variations in hypervector size. Likewise, as depicted in Table 3, the number of training samples also does not influence ${\sigma}^{*}_{\text{b}}$. Thus, we deduce that ${\sigma}^{*}_{\text{b}}$ is unrelated of these DP-HD parameters and the scale of training data. It is solely contingent upon the distribution characteristics of the feature vectors, such as the proximity between class centers and the spread of samples around these centers.
@@
Subsequently, we analyzed our dataset to determine the optimal ${\sigma}^{*}_{\text{b}}$ for our application. The core objective of our research was to find the best compromise between robust privacy protection and high operational efficiency. Figure 8 illustrates how the DP-HD model performs under different $\epsilon$ levels with both exclusive and inclusive encoding strategies. Our efforts centered around finding a ${\sigma}^{*}_{\text{b}}$ that achieves high accuracy, demonstrating superior performance, while preserving a low $\epsilon$, which reflects stronger privacy. In previous studies of HD, ${\sigma}_{\text{b}}$ is fixed at one, which we refer to as BaselineHD, the accuracy remains below 15% across all $\epsilon$ values, underscoring the significance of adapting ${\sigma}_{\text{b}}$ to balance privacy and performance effectively. In DP-HD, we selected ${\sigma}^{*}_{\text{b}}$, optimizing both accuracy and privacy. At this configuration, the model achieves the highest accuracy of 97.6% with the least amount of noise, corresponding to the highest $\epsilon$ of 10. As $\epsilon$ decreases, privacy increases, but at the expense of accuracy. For instance, when $\epsilon$ falls to 0.2, accuracy declines below 70%. Conversely, when $\epsilon$ exceeds 0.5, accuracy remains above 90%. It is clear that for any level of privacy, as determined by choosing a specific value of $\epsilon$, the ${\sigma}^{*}_{\text{b}}$ will be changed. This happens since there is a randomness in generating the random basis vectors that effect the model’s performance. However, it is apparent for all privacy $\epsilon$ values, the ${\sigma}^{*}_{\text{b}}$ is close to 0.2. Therefore we chose 0.2 as ${\sigma}^{*}_{\text{b}}$ to assign our dataset.
The uncertainty in DP-HD’s output can be assessed by examining how much accuracy changes at a given privacy level when $\sigma_{\text{b}}$ deviates from the optimal value, $\sigma^{*}_{\text{b}}$. Specifically, we identify the range of $\sigma_{\text{b}}$ values around $\sigma^{*}_{\text{b}}$ that still allow accuracy to remain above 90%, as illustrated in Figure 8. For instance, at $\epsilon=10$, selecting $\sigma_{\text{b}}$ between 0.03 and 0.35 ensures accuracy exceeds 90%, which corresponds to a substantial 160% variance from $\sigma^{*}_{\text{b}}=0.2$. This demonstrates that the DP-HD framework maintains a stable balance between accuracy and privacy even with a large range of $\sigma_{\text{b}}$ deviations.
As $\epsilon$ decreases, indicating a stronger privacy requirement, the range of acceptable $\sigma_{\text{b}}$ values narrows slightly but remains robust. For example, at $\epsilon=5$ and $\epsilon=2$, $\sigma_{\text{b}}$ can still vary between 0.06 and 0.33, offering 135% flexibility around $\sigma^{*}_{\text{b}}$ while sustaining accuracy above 90%. At $\epsilon=1$, this range narrows further to 0.08–0.28, and at the high privacy level of $\epsilon=0.5$, $\sigma_{\text{b}}$ can vary between 0.11 and 0.2, maintaining 45% flexibility in $\sigma_{\text{b}}$ selection. These findings indicate that DP-HD’s output uncertainty is minimal; by choosing an appropriate range for $\sigma_{\text{b}}$, high accuracy is maintained across various privacy levels, underscoring the framework’s reliability in balancing accuracy and privacy.
@@
5.1.2 Benchmark Comparison
In this analysis, we benchmarked our newly developed DP-HD model against established models in the field of image classification. This comparison includes an evaluation alongside models such as ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2. To implement DP, we utilized Opacus, a freely available open-source library designed for PyTorch that facilitates training deep learning models with DP mechanisms. Table 4 details the accuracies post-DP application, revealing that DP-HD sustains a minor accuracy reduction compared to significant drops in other models. With a stringent privacy level, represented by an $\epsilon$ value of 0.6, DP-HD maintains an accuracy of 91.04%, whereas other models plummet below 43%, showcasing DP-HD’s robustness against high noise levels as it distributes memory evenly across all dimensions of the hyperspace. As $\epsilon$ increases to 0.8, reducing the noise, DP-HD’s accuracy climbs to 93.34%, significantly surpassing other models which remain under 54%. This performance gap widens further when $\epsilon$ is adjusted to 1, with DP-HD reaching an accuracy of 94.43%, and the closest competitor, DenseNet201, achieves only 69.13%, still 25.3% lower than DP-HD. As $\epsilon$ increases further to 2, the accuracy of DP-HD is 96.30%, while the other models’ accuracies—75.22% for ResNet50, 82.48% for GoogLeNet, 58.20% for AlexNet, 77.70% for DenseNet201, and 57.20% for EfficientNet B2—illustrate a substantial disparity, with GoogLeNet being the closest yet still 13.82% less accurate than DP-HD. Even at an $\epsilon$ of 3, where privacy levels are lower and noise reduction is significant, DP-HD’s accuracy remains high at 96.66%, outperforming ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2, which exhibit accuracies of 81.18%, 83.73%, 56.27%, 86.89%, and 56.46% respectively. This demonstrates DP-HD’s superior capability to balance privacy and performance compared to other prevalent ML models in the field.